OpenStudy (anonymous):

Can someone help me create a graph for f(x)=-3cot((x/2)-(pi/3))

OpenStudy (amorfide):

you could change cot(x/2) to 1/(tan(x/2)) that should help you, then just substitute values

OpenStudy (amorfide):

or better yet, change cot(x/2) to \[\frac{ \cos(x/2) }{ \sin(x/2)}\]

OpenStudy (amorfide):

then sub in x values